OpenStudy (anonymous):

College Algebra: Functions and Graphs: Find x- and y-intercepts: -1/2x+y-4=0 Stuck on finding y-intercept, My work so far: AX+BY+C=0 X-int; y=0 -1/2X+0-4=0 -1/2x=4 x=-2, x-int: (-2,0) *** y-int; x=0 -1/2(0)+y-4=0 0+y-4=0 y=4; y-int: (0,4) <--- answer key says its (-8,0) though... how??

OpenStudy (gorv):

to find x intercept put y=0

OpenStudy (gorv):

to find y intercept put x=0

TheSmartOne (thesmartone):

@gorv You forgot to say Welcome to OS!!

OpenStudy (anonymous):

Okay...y-int; x=0 -1/2(0)+y-4=0 0+y-4=0 y=4; y-int: (0,4) <--- answer key says its (-8,0) though... how??